Use induction and Pascals identity to prove that () -2 nzo и n where

Using Induction and Pascal's Identity

Using Mathematical Induction
Answer #1

Claim & () = 2. By induction , n=1 1 basis. step Ź (j) = 2 = 2 LHS j 20 R.HiS = (0) + (1) = 1+1=2. let it is true for na